# Probing arbitrary polarized photon pairs undergoing double Compton scatterings by a dedicated MC simulator validated with experimental data

### v1

@@ Abstract
Quantum correlations in the degree of polarization of freedom of the two-photon system have been extensively studied and form our current understanding of the quantum nature of our world. Most of the studies are concentrated on the low-energy (optical) photon pairs, for which efficient polarization measurement devices exist. However, for high-energetic (MeV) pairs of photons, e.g. produced in the decay of positronium atoms, no polarizers are available. Partial information about the polarization degree of freedom can be extracted by exploiting the measurements of photon pairs that undergo double Compton scattering. We present a Geant4-based Monte Carlo Vienna-Warsaw model capable of simulating any initial polarization state of bipartite photons. This puts us in a position to derive the behavior of the experimental observable, the angular difference $\Delta\hat{\Phi}$ formed by the two scattering planes. We validate our Vienna-Warsaw simulator with the high-statistics experimental sample –based on a total of $3\times 10^{5}$ event candidates — of two-photon pairs measured with the J-PET Big Barrel detector. We deduce the value of the squared visibility (interference contrast) encoding the polarization in the angle difference of the two scattering planes, $\Delta\hat{\Phi}$. The simulated spectra are in good agreement with the experimental correlation spectra and behave as predicted by theory.
Keywords:
quantum correlations, Compton scattering, positron emission tomography, bipartite system
@@ 1 Introduction
Polarisation measurements are widely used to show the very working of quantum mechanics and, in particular, quantum correlations of many-photon systems are investigated, providing a way to encode information for quantum information theoretic tasks such as e.g. quantum computing Madsen et al. (2022), quantum machine learning Hiesmayr (2022); Biamonte et al. (2017); Wan et al. (2017) or quantum cryptography Bennett and Brassard (2014); Ekert (1991); Schauer et al. (2010); Bennett et al. (1992); Cao et al. (2020). Most of the studies so far have been concentrated on optical photon systems, in which the energies of the photons are in the range of a few eV. However, recent investigations based on the estimation of the polarization for the MeV photon pairs originating from the decay of the positronium atoms or the direct electron-positron annihilation have regained attention. Particularly, the interest is enhanced by the proposal to exploit it in medical applications, i.e. in the context of positron emission tomography (PET)  McNamara et al. (2014); Toghyani et al. (2016); Moskal et al. (2018); Bass et al. ; Watts (2024). The main idea is to use the correlation observable - the relative angle of the two-photon scattering planes $\Delta\hat{\Phi}$ to distinguish the genuine, undisturbed photon pairs from the noisy events formed by the uncorrelated (random) photons or photons scattered in the patient’s body. The reduction of this background could improve the final PET image contrast Kožuljević et al. (2021). Indeed, recently several pilot measurements were performed Makek et al. (2020); Watts et al. (2021); Abdurashitov et al. (2022). Interestingly, the recent results dedicated to the analysis of the scattered photons  Ivashkin et al. (2023); Parashari et al. (2024); Bordes et al. (2024) seem to contradict the previous conclusions Watts et al. (2021). The applicability of the proposed methodology in practical scanners remains, therefore, an open question, nevertheless, further investigations are ongoing Kim et al. (2023); Shoop et al. (2024); Das et al. (2024); Kožuljević et al. (2024); Watts (2024); Moulin et al. (2024); Moskal et al. (2025).
@@
In this work, we present studies of photon correlations probed by the double Compton scattering events based on the theory published in Ref.  Hiesmayr and Moskal (2019). The investigations are based on the Monte Carlo (MC) simulations utilising the developed ‘Vienna-Warsaw model’(VW) by implementing it into the Geant4 simulation software. Namely, our VW simulation model allows computing the expected correlation of any initial bipartite state being either separable or entangled, mixed or pure. MC-based predictions are validated with the experimental spectra obtained based on the analysis of the high-statistics experimental sample measured by the Jagiellonian Positron Emission Tomograph (J-PET) in 2020 Niedźwiecki et al. (2017); Dulski et al. (2021)111 The experimental input dataset used in this article is overlapping with the one used in Moskal et al. (2025). The first version of the current article was finalised on 16.08.2023. The submission was postponed at the request of the J-PET group leader P. Moskal, who wanted to prioritize Moskal et al. (2025)..
@@ 2.1 Klein-Nishina formula for single photons detected by plastic scintillators
{dmath}dσfi
= r022K^2(k,~Θ)(K^-1(k,~Θ)+K(k,~Θ)-2+4|ε_f^*⋅ε_i |^2) where $i/f$ refer to the initial/final polarised states described by $\varepsilon$, $r_{0}$ is the classical electron radius, and $k$ is the energy of the incoming photon (see Fig. 1), and
@@
$\frac{d}{d\Omega}\sigma_{\rho}(\theta,\phi,\theta_{a},\phi_{a},k)=\frac{r_{0}^{2}}{2}\;K^{2}(k,\tilde{\Theta})\;\sum_{i=1}^{2}\text{Tr}\left[\mathcal{K}_{i}\rho\mathcal{K}^{\dagger}\right]$ (3)
@@
Any initial state $\rho$ can be written by an orthonormal basis ($\langle H|V\rangle=0,\langle H|H\rangle=\langle V|V\rangle=1$) {dmath} ρ= ρ_HH|H⟩⟨H|+ρ_HV|H⟩⟨V|+ρ_VH|V⟩⟨H| +ρ_VV|V⟩⟨V| where $\rho_{HH},\rho_{VV}$ are positive numbers that add up to $\rho_{HH}+\rho_{VV}=1$ and $\rho_{HV}=\rho_{VH}^{*}$ are complex numbers and generally the positivity $\rho\geq 0$ has to hold. Given this basis choice, the Kraus operators decomposed into Pauli matrices $\hat{\sigma}_{i}$ are defined by
@@
{dmath}
K_2 = (cosθacosθ+1)cos(ϕ-ϕa)+sinθasinθ2 1_2 + (cosθacosθ-1)cos(ϕ-ϕa)+sinθasinθ2 ^σ_3 + i sin(ϕ-ϕa)(cosθ-cosθa)2 ^σ_1 + sin(ϕ-ϕa)(cosθ+cosθa)2 ^σ_2 . The direction of motion and a chosen coordinate system are connected via the Compton scattering angle {dmath} cos~Θ=^k⋅^k_a = cos(θ_a-θ) + (cos(ϕ-ϕ_a)-1)sinθ_asinθ .
@@
$\mathcal{F}(\tilde{\Theta},k)=K^{2}(k,\tilde{\Theta})\biggl{(}K^{-1}(k,\tilde{\Theta})+K(k,\tilde{\Theta})-\sin^{2}\tilde{\Theta}\biggr{)}$ (6)
@@
Figure 2: (a) The envelope function $\mathcal{F}$ for the given incoming photon energy $k$ and the scattering angle $\tilde{\Theta}$ defines the maximum amplitude of the visible oscillations for polarized photons. The line colors represent different energies: black $k=0.1$, red $k=1(511~{}\text{keV})$, green $k=2$ and blue $k=5$. (b) The visibility or interference contrast is the component that reduces the oscillation patterns depending on the incoming photon energy $k$ and the scattering angle $\tilde{\Theta}$. In particular, if the visibility is 00, then no oscillation in $\Phi$ can be observed. This is the case for all energies if we consider small-angle scatterings or backwards scattering.
@@
In particular, if we assume the source is producing an unpolarised state, i.e. $\rho_{HH}=\rho_{VV}=\frac{1}{2}$ and $\rho_{HV}=\rho_{VH}=0$, equation (Eq. 5) predicts that no $\Phi$ dependence can be observed, whereas if the source produces e.g. a pure fully polarised state, e.g. |$\ket{\text{H}}$, it predicts an oscillatory behaviour with $\Phi$. Surprisingly at first glance, the formula also predicts a vanishing oscillation for a pure $|45\rangle$-polarized state. This is due to the summation over all final states, assuming the electrons in the plastic to be totally unpolarized.
@@ 2.2 The Klein-Nishina formula for multiphotonic Compton scattering events
where $a_{i},b_{i},c_{ij}$ are real numbers. Considering only locally maximally mixed states, i.e. setting $a_{i}=b_{i}=0$, the remaining state space corresponds to a three-dimensional tetrahedron, a magic simplex  Baumgartner et al. (2007); Hiesmayr (2021), where the maximally entangled states are represented by vertices. All those states are equal in the property that there is only information encoded in the whole system, but not in any subsystems (individual photons). However, note that not all polarization states are possible if the bosonic nature of photons, i.e. that the total wave function has to be symmetric under particle exchange, is taken into account (for more details see the appendix in Ref.  Hiesmayr and Moskal (2019)).
@@ 3 Interpretation of the Compton scattering as an estimator of polarisation
[…]
Next, we compute the angle region around $\Phi_{max}=\frac{\pi}{2},\frac{3\pi}{2}$ such that it covers $68\%$ of the events and we interpret the half of this region as one standard deviation. This gives a value of about $42^{\circ}$. Therefore, if we considered a source of photons with the same $|H\rangle$ initial polarization state, then in $68\%$ of the cases the angle between the scattering planes and the initial polarization direction would lay in a range not wider than $[48^{\circ},132^{\circ}]$ or $[228^{\circ},312^{\circ}]$.
@@ 4.1 VW Model applied to different initial bipartite states for ideal detectors
Figure 6: The distributions of azimuthal angles $\phi_{a}$ versus $\phi_{b}$ of primary photon pairs under the assumption of perfect detection efficiency for four different initial quantum states. The difference between the histograms in (a,Eq. 12) and (b,Eq. 13) is a shift of $90^{\circ}$ corresponding to the definition of |$\ket{H}$ and |$\ket{V}$ assuming the classic mechanics model. Histogram (c,Eq. 14 and Eq. 4.1) can be obtained in two ways: as a statistical mixture of state |$\ket{HV}$ and |$\ket{VH}$ or for the entangled state |$\ket{\psi^{+}}$ in both cases, we assumed the quantum mechanics model. We assumed here that both annihilation photons propagate along the $z$-axis.
@@
Consequently, adding the weighted sum of the two individual simulations (Fig. 6(a)/ Fig. 6(b)) leads to Fig. 6(c). However, assuming $|\psi^{+}\rangle$ leads also to Fig. 6(c), but here each event has the same individual statistics behind each of both cases |$\ket{HV}$ and |$\ket{VH}$ in the simulation, in strong contrast to the previous case. One observes the strips of destructive and constructive interference.
@@ 4.2 Comparison of ‘VW Simulation Model’ with current simulators
Figure 7: Distribution of the $\Delta\hat{\Phi}$ observable generated by the VW model for (a) all Compton scattering angles and (b) restricted scattering angles to the circle area with the radius of $30^{\circ}$ and the centre placed at the maximal visibility angles for 511 keV photons $\tilde{\Theta}_{\rm{max}}=\tilde{\Theta}_{a}=\tilde{\Theta}_{b}=81.66^{\circ}$. The (black triangles) graph corresponds to the result assuming an initial maximally entangled state |$\ket{\psi^{+}}$, the (red down triangles) graph assumes the classical mechanics model of polarised photons, and the (blue diamonds) graph corresponds to the initial unpolarized state. The last two scenarios agree with the results of the Polarized Livermore Model and Standard Livermore Model, respectively. Restricting the Compton angle region leads to higher visibility but lowers the overall statistics.
In the Livermore Polarized Model the linear polarization of the individual photons can be chosen to be orthogonal to each other. This effectively leads to the $\Delta\hat{\Phi}$ distribution which would correspond to the simulation of a separable quantum state of type |$\ket{\text{HV}}$. Indeed, the same shape is reproduced by the VW model with the polarization state set to |$\ket{\text{HV}}$ (see Fig. 7 red down triangles.).
@@ 5.1 Detection system and positronium source
The study is performed based on a subsample of the data collected in 2020 by the J-PET Big Barrel detector Niedźwiecki et al. (2017); Dulski et al. (2021). It consists of three scintillator layers oriented cylindrically. Each scintillator has $19$ mm $x$-length, $7$ mm $y$-length and $500$ mm $z$-length. The first layer has $48$ scintillators and a radius of $425$ mm, the second layer has $48$ scintillators shifted relative to the first layer with $3.75^{\circ}$ and a radius of $467.5$ mm, the third layer has $96$ scintillators shifted relative to the first layer with $1.875^{\circ}$ and a radius of $575$ mm. The detector readout system consists of the Hamamatsu R9800 vacuum photomultipliers connected from both ends to each EJ-230 scintillator strip, which transforms the generated light into electric signals. The signals are probed at four thresholds and stored by the dedicated triggerless data acquisition system (DAQ) Korcyl et al. (2018).
@@ 5.3 MC samples
|$\ket{\psi^{+}(\pm 45^{\circ})}=\frac{1}{\sqrt{2}}\left(\ket{+45^{\circ}-45^{\circ}}+\ket{-45^{\circ}+45^{\circ}}\right)$
@@
The simulated MC samples consisted of $2\cdot 10^{12}$ photon pairs, generated using the experimental setup corresponding to the J-PET Big Barrel detector. During the post-processing, the registered positions, times and energies were smeared using the phenomenological parameterizations modelling the experimental resolutions of the front-end electronics and the DAQ system. The energy resolution dependence is parameterised as a $\frac{\sigma(E)}{E}=\frac{0.06}{\sqrt{\Delta E[MeV]}}$ fraction. The simulated photon registration time is smeared, event by event, by replacing the event registration time $t_{r}$ by the value obtained from the normal distribution $N(t_{r},\sigma_{t}$), where $\sigma_{t}=0.22\sqrt{\frac{340}{\Delta E[keV]}}~{}\text{ns}$ corresponds to the temporal resolution. Analogically, the registration position along the scintillator strips ($z$ position) is smeared, event by event, by replacing the registered photon position $z$ by the value obtained from the normal distribution $N(z,\sigma_{z})$, where $\sigma_{z}=2.5~{}\text{cm}$ corresponds to the positional uncertainty along the scintillator strip of the scanner. To mimic the measurement conditions and the TOT selection, we accepted only the events which consisted of four hits with the deposited energy lying in the range from 70 keV to 340.6 keV. The sample was subsequently filtered by applying the same selection procedure as in the case of the experimental data. Out of all the generated events, the $4.18\times 10^{7}$ candidates fulfilled 4-hit conditions, and about $2.24\times 10^{5}$ remained after applying the final selection procedure.

### v2

@@ Abstract
Quantum correlations in the polarization degrees of freedom of the two-photon system have been extensively studied and form our current understanding of the quantum nature of our world. Most of the studies are concentrated on the low-energy (optical) photon pairs, for which efficient polarization measurement devices exist. However, for high-energetic (MeV) pairs of photons, e.g. produced in the decay of positronium atoms, no polarizers are available. Partial information about the polarization degree of freedom can be extracted by exploiting the measurements of photon pairs that undergo double Compton scattering. We present a Geant4-based Monte Carlo Vienna-Warsaw model capable of simulating any initial polarization state of bipartite photons. This puts us in a position to derive the behavior of the experimental observable, the angular difference $\Delta\hat{\Phi}$ formed by the two scattering planes. We validate our Vienna-Warsaw simulator with the high-statistics experimental sample –based on a total of $3\times 10^{5}$ event candidates — of two-photon pairs measured with the J-PET Big Barrel detector. We deduce the value of the squared visibility (interference contrast) encoding the polarization in the angle difference of the two scattering planes, $\Delta\hat{\Phi}$. The simulated spectra are in good agreement with the experimental correlation spectra and behave as predicted by theory.
@@ 1 Introduction
Polarisation measurements are widely used to show the very working of quantum mechanics and, in particular, quantum correlations of many-photon systems are investigated, providing a way to encode information for quantum information theoretic tasks such as e.g. quantum computing Madsen et al. (2022), quantum machine learning Hiesmayr (2022); Biamonte et al. (2017); Wan et al. (2017) or quantum cryptography Bennett and Brassard (2014); Ekert (1991); Schauer et al. (2010); Bennett et al. (1992); Cao et al. (2020). Most of the studies so far have been concentrated on optical photon systems, in which the energies of the photons are in the range of a few eV. However, recent investigations based on the estimation of the polarization for the MeV photon pairs originating from the decay of the positronium atoms or the direct electron-positron annihilation have regained attention. Particularly, the interest is enhanced by the proposal to exploit it in medical applications, i.e. in the context of positron emission tomography (PET)  McNamara et al. (2014); Toghyani et al. (2016); Moskal et al. (2018); Bass et al. ; Watts (2024). The main idea is to use the correlation observable - the relative angle of the two-photon scattering planes $\Delta\hat{\Phi} -$ to distinguish the genuine, undisturbed photon pairs from the noisy events formed by the uncorrelated (random) photons or photons scattered in the patient’s body. The reduction of this background could improve the final PET image contrast Kožuljević et al. (2021). Indeed, recently several pilot measurements were performed Makek et al. (2020); Watts et al. (2021); Abdurashitov et al. (2022). Interestingly, the recent results dedicated to the analysis of the scattered photons  Ivashkin et al. (2023); Parashari et al. (2024); Bordes et al. (2024) seem to contradict the previous conclusions Watts et al. (2021). The applicability of the proposed methodology in practical scanners remains, therefore, an open question, nevertheless, further investigations are ongoing Kim et al. (2023); Shoop et al. (2024); Das et al. (2024); Kožuljević et al. (2024); Watts (2024); Moulin et al. (2024); Moskal et al. (2025).
@@
In this work, we present studies of photon correlations probed by the double Compton scattering events based on the theory published in Ref.  Hiesmayr and Moskal (2019). The investigations are based on the Monte Carlo (MC) simulations utilising the developed ‘Vienna-Warsaw model’(VW) by implementing it into the Geant4 simulation software. Namely, our VW simulation model allows computing the expected correlation of any initial bipartite state being either separable or entangled, mixed or pure. MC-based predictions are validated with the experimental spectra obtained based on the analysis of the high-statistics experimental sample measured by the Jagiellonian Positron Emission Tomograph (J-PET) in 2020 Niedźwiecki et al. (2017); Dulski et al. (2021) [^1] The experimental input dataset used in this article is overlapping with the one used in Moskal et al. (2025). The first version of the current article was finalised on 16.08.2023. The submission was postponed at the request of the J-PET group leader P. Moskal, who wanted to prioritize Moskal et al. (2025)..
@@ 2.1 Klein-Nishina formula for single photons detected by plastic scintillators
$\frac{d\sigma_{fi}}{d\Omega}=\frac{r_{0}^{2}}{2}K^{2}(k,\tilde{\Theta})\left(K^{-1}(k,\tilde{\Theta})+K(k,\tilde{\Theta})-2+4\lvert{\boldsymbol{\varepsilon}_{f}}^{*}\cdot\boldsymbol{\varepsilon}_{i}\rvert^{2}\right)$ (1)
where $i/f$ refer to the initial/final polarised states described by $\varepsilon$, $r_{0}$ is the classical electron radius, and $k$ is the energy of the incoming photon expressed in $\frac{\text{keV}}{m_{e}c^{2}}$ (see Fig. 1), and
@@
$\frac{d}{d\Omega}\sigma_{\rho}(\theta,\phi,\theta_{a},\phi_{a},k)=\frac{r_{0}^{2}}{2}\;K^{2}(k,\tilde{\Theta})\;\sum_{i=1}^{2}\text{Tr}\left[\mathcal{K}_{i}\rho\mathcal{K}_{i}^{\dagger}\right]$ (4)
@@
Any initial state $\rho$ can be written by an orthonormal basis ($\langle H|V\rangle=0,\langle H|H\rangle=\langle V|V\rangle=1$)
[…]
where $\rho_{HH},\rho_{VV}$ are positive numbers that add up to $\rho_{HH}+\rho_{VV}=1$ and $\rho_{HV}=\rho_{VH}^{*}$ are complex numbers and generally the positivity $\rho\geq 0$ has to hold. Given this basis choice, the Kraus operators decomposed into Pauli matrices $\hat{\sigma}_{i}$ are defined by
@@
$\mathcal{K}_{2}=\frac{(\cos\theta_{a}\cos\theta+1)\cos(\phi-\phi_{a})+\sin\theta_{a}\sin\theta}{2}\;\mathbbm{1}_{2}+\frac{(\cos\theta_{a}\cos\theta-1)\cos(\phi-\phi_{a})+\sin\theta_{a}\sin\theta}{2}\;\hat{\sigma}_{3}+\frac{i\sin(\phi-\phi_{a})(\cos\theta-\cos\theta_{a})}{2}\;\hat{\sigma}_{1}+\frac{\sin(\phi-\phi_{a})(\cos\theta+\cos\theta_{a})}{2}\;\hat{\sigma}_{2}\;.$ (7)
The direction of motion and a chosen coordinate system are connected via the Compton scattering angle
$\cos\tilde{\Theta}=\hat{\bf{k}}\cdot\hat{\bf{k_{a}}}=\cos(\theta_{a}-\theta)+\left(\cos(\phi-\phi_{a})-1\right)\sin\theta_{a}\sin\theta\;.$ (8)
@@
[…]
@@
Figure 2: (a) The envelope function $\mathcal{F}$ for the given incoming photon energy $k$ and the scattering angle $\tilde{\Theta}$ defines the maximum amplitude of the visible oscillations for polarized photons. The line colors represent different energies: black $k=0.1$, red $k=1(511~\text{keV})$, green $k=2$ and blue $k=5$. (b) The visibility or interference contrast is the component that reduces the oscillation patterns depending on the incoming photon energy $k$ and the scattering angle $\tilde{\Theta}$. In particular, if the visibility is 0, then no oscillation in $\Phi$ can be observed. This is the case for all energies if we consider small-angle scatterings or backwards scattering.
@@
[…]
@@ 2.2 The Klein-Nishina formula for multiphotonic Compton scattering events
where $a_{i},b_{i},c_{ij}$ are real numbers. Considering only locally maximally mixed states, i.e. setting $a_{i}=b_{i}=0$, the remaining state space corresponds to a three-dimensional tetrahedron, a magic simplex  Baumgartner et al. (2007); Hiesmayr (2021), where the maximally entangled states are represented by vertices. All those states are equal in the property that there is only information encoded in the whole system, but not in any subsystems (individual photons). However, note that not all polarization states are possible if the bosonic nature of photons, i.e. that the total wave function has to be symmetric under particle exchange, is taken into account (for more details see Ref.  Hiesmayr and Moskal (2019)).
@@ 3 Interpretation of the Compton scattering as an estimator of polarisation
$f(\Phi|\tilde{\Theta}_{\rm{max}})$ $=$ $\frac{\frac{d\sigma_{\rho}(\tilde{\Theta}_{\rm{max}},\Phi)}{d\Omega}}{\int_{0}^{2\pi}\;\frac{d\sigma_{\rho}(\tilde{\Theta}_{\rm{max}},\Phi)}{d\Omega}\;d\Phi}$ (14)
Next, we compute the angle region around $\Phi_{max}=\frac{\pi}{2},\frac{3\pi}{2}$ such that it covers $68\%$ of the events and we interpret the half of this region as one standard deviation. This gives a value of about $42^{\circ}$. Therefore, if we considered a source of photons with the same $|H\rangle$ initial polarization state, then in $68\%$ of the cases the angle between the scattering planes and the initial polarization direction would lay in a range not wider than $48^{\circ},132^{\circ}]$ or $228^{\circ},312^{\circ}]$.
@@ 4.1 VW Model applied to different initial bipartite states for ideal detectors
[…]
@@
[…]
@@ 4.2 Comparison of ‘VW Simulation Model’ with current simulators
[…]
@@ 5.1 Detection system and positronium source
The study is performed based on a subsample of the data collected in 2020 by the J-PET Big Barrel detector Niedźwiecki et al. (2017); Dulski et al. (2021). It consists of three scintillator layers oriented cylindrically. Each scintillator has $19$ mm $x$-length, 77 mm $y$-length and $500$ mm $z$-length. The first layer has $48$ scintillators and a radius of $425$ mm, the second layer has $48$ scintillators shifted relative to the first layer with $3.75^{\circ}$ and a radius of $467.5$ mm, the third layer has $96$ scintillators shifted relative to the first layer with $1.875^{\circ}$ and a radius of $575$ mm. The detector readout system consists of the Hamamatsu R9800 vacuum photomultipliers connected from both ends to each EJ-230 scintillator strip, which transforms the generated light into electric signals. The signals are probed at four thresholds and stored by the dedicated triggerless data acquisition system (DAQ) Korcyl et al. (2018). The spatial resolution along the scintillator strip is equal to $\sigma=2.5~\text{cm}$, while the Time-of-Flight resolution corresponds to about 220 ps  Niedźwiecki et al. (2017).
@@ 5.3 MC samples
[…]
@@
The simulated MC samples consisted of $2\cdot 10^{12}$ photon pairs, generated using the experimental setup corresponding to the J-PET Big Barrel detector. During the post-processing, the registered positions, times and energies were smeared using the phenomenological parameterizations modelling the experimental resolutions of the front-end electronics and the DAQ system. The energy resolution dependence is parameterised as a $\frac{\sigma(E)}{E}=\frac{0.06}{\sqrt{\Delta E[MeV]}}$ fraction. The simulated photon registration time is smeared, event by event, by replacing the event registration time $t_{r}$ by the value obtained from the normal distribution $N(t_{r},\sigma_{t}$), where $\sigma_{t}=0.22\sqrt{\frac{340}{\Delta E[keV]}}~\text{ns}$ corresponds to the temporal resolution. Analogically, the registration position along the scintillator strips ($z$ position) is smeared, event by event, by replacing the registered photon position $z$ by the value obtained from the normal distribution $N(z,\sigma_{z})$, where $\sigma_{z}=2.5~\text{cm}$ corresponds to the positional uncertainty along the scintillator strip of the scanner. Due to limitations of the front-end electronics, some events would involve 511 keV photons undergoing multiple interactions within the same scintillator strip. In such cases, all interactions can be recorded as a single signal. This effect is partially accounted for in our simulations by merging multiple scattering within the experimental resolution. However, it should be emphasised that it remains unclear whether the Geant4 LivermorePolarized model accurately describes the final polarisation state of photons after multiple scatterings in this context. A more precise treatment would require the implementation of a dedicated multiple scattering model as proposed, for example, in Hiesmayr et al. (2024); Caradonna (2024). To mimic the measurement conditions and the TOT selection, we accepted only the events which consisted of four hits with the deposited energy lying in the range from 70 keV to 340.6 keV. The sample was subsequently filtered by applying the same selection procedure as in the case of the experimental data. Out of all the generated events, the $4.18\times 10^{7}$ candidates fulfilled 4-hit conditions, and about $2.24\times 10^{5}$ remained after applying the final selection procedure.